OpenStudy (izzyrawrz):

1. Solve each equation. Show all your work. Round your answers to four decimal places. a. 6^(3x) = 104 b. log 2 + log(4x – 1) = 3 I really need help here with this. Danke!

OpenStudy (welshfella):

a take logs of both sides log 6^3x = log 104 3x log g = log 104 solve for x

OpenStudy (welshfella):

* that is log 6 not log g - typo

OpenStudy (welshfella):

3x log 6 = log 104 3x = log 104 / log 6 x = log 104 / 3 log 6 grab your calculator and you've got it.

OpenStudy (welshfella):

did you follow the solution I used the rule which says log a^b = b log a

OpenStudy (welshfella):

b . log 2 + log (4x -1) = 3 log 2(4x - 1) = 3 we are using logs to the base 10 so 2(4x -1) = 10^3

OpenStudy (welshfella):

now that's easy to solve.

OpenStudy (welshfella):

are you there?

OpenStudy (izzyrawrz):

Danke, but let see if i did it correctly the answer for part (a) is .86 and part (b) would be 125.125 am I right? P.S. I was doing something else sorry. @welshfella

OpenStudy (loser66):

yes for a

OpenStudy (loser66):

yes for b

OpenStudy (izzyrawrz):

Danke! @Loser66